Circular No. 7840

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

γ-RAY TRANSIENT S. Golenetskii, R. Aptekar, E. Mazets, and D. Frederiks, Ioffe Insti-

tute, St. Petersberg, on behalf of Konus-Wind γ-ray-burst (GRB) team; T. Cline, Goddard Space Flight Center, NASA, on behalf of the Konus-Wind and Ulysses GRB teams; and K. Hurley, Space Sciences Laboratory, University of California at Berkeley, on behalf of the Ulysses GRB team, report that Konus-Wind and Ulysses observed an exceptional γ-ray transient from a source with high north-ecliptic latitude of duration > 10 000 s, beginning ∼ Feb. 24.625 UT. Triangulation gives an annulus centered at α = 17^{h}21^{m}46^{s}.3, δ = +78^{o}12′29′′ (equinox 2000.0), with radius 46^{o}.138 ± 1^{o}.215. This localization is inconsistent with a solar origin, and the error box excludes all known soft-γ-ray repeaters (SGR), but it does include Cygnus X-1. In the past seven years, at least two similar transients have been observed by Konus and Ulysses (cf. Mazets et al. 1995, AIP Conf. Proc. 384, 492). Konus-Ulysses triangulation of the two 1995 events gives the following annulus centers and radii: Jan. 10.271–10.368, α = 21^{h}28^{m}32^{s}.4, δ = −45^{o}45′32′′, 82^{o}.981 ± 1^{o}.574; Mar. 25.258–25.302, α = 23^{h}15^{m}40^{s}.1, δ = +0^{o}56′31′′, 58^{o}.963 ± 1^{o}.782. These localizations are similarly inconsistent with a solar or SGR origin. The three annuli intersect to form an ≈ 8-deg^{2} error box that includes Cyg X-1. M. Briggs (private comm.) attributed the Mar. 25 event to Cyg X-1. It is concluded that all three outbursts originated from Cyg X-1 while in an exceptionally high state. Fluences (units of 10^{−4} erg cm^{−2}) and peak fluxes (10^{−7} erg cm^{−2} s^{−1}) of these events above 15 keV, and corresponding energies (10^{41} erg) and luminosities (10^{38} erg/s), are: 1995 Jan. 10, 8.4, 2, 6.25, 1.5; Mar. 25, 3.0, 3, 2.2, 2.2; 2002 Feb. 24, 1.5, 1.5, 1.1, 1.1. These fluxes are an order of magnitude higher than fluxes measured by RXTE during a peculiar flaring episode of Cyg X-1 one year ago (Cui et al. 2002, Ap.J. 564, L78). This is the first time that the Interplanetary Network has been used to localize a cosmic transient source that is not an SGR or a GRB.

SUPERNOVA 2002bh IN UGC 5286 W. M. Wood-Vasey, University of California at Berkeley and Lawrence

Berkeley National Laboratory, reports that pre-discovery CCD images of SN 2002bh (cf. IAUC 7837) taken with the Oschin 1.22-m telescope at Palomar on Feb. 16.4 UT show the new object at unfiltered mag 17.6.

2002 March 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green